### v2

# Thomason condition for regular algebraic stacks

Abstract.
We show that the Thomason condition holds for any concentrated regular Noetherian algebraic stack with quasi-finite and locally separated diagonal. As a consequence, the derived category of quasi-coherent sheaves on such an algebraic stack is singly compactly generated. This extends results of Hall–Rydh from separated to locally separated diagonal, under additional regularity hypotheses.
Key words and phrases:
Algebraic stacks, perfect complexes, Thomason condition, derived categories, compact objects
2020 Mathematics Subject Classification:
14A30 (primary), 14D23, 14F08, 18G80
1. Introduction
1.1. What is known
Compact generation of derived categories of quasi-coherent sheaves on algebraic stacks plays a central role in modern algebraic geometry. In many cases of interest, the triangulated category $D_{\operatorname{qc}}$ is singly compactly generated. That is, there exists a single compact object whose shifts, cones, and small coproducts generate $D_{\operatorname{qc}}$.
For quasi-compact quasi-separated schemes, $D_{\operatorname{qc}}$ being singly compactly generated was established by Bondal and Van den Bergh [BV03, Theorem 3.1.1]. Extensions to various classes of Deligne–Mumford and algebraic stacks—typically in characteristic zero or under resolution property assumptions—were obtained in [Toë12, Corollary 5.2], [Kri09, proof of Proposition 5.5], and [BZFN10, §3.3].
A recent advance was made by Hall and Rydh, who proved that $D_{\operatorname{qc}}$ is singly compactly generated for any quasi-compact quasi-separated algebraic stack with quasi-finite and separated diagonal [HR17a, Theorem A]. In that work, Hall and Rydh introduced the $\beta$-Thomason condition, where $\beta$ is a cardinal. This requires not only compact generation of $D_{\operatorname{qc}}$ by a collection of cardinality $\leq\beta$, but also the existence of perfect complexes with prescribed support. Note that the $1$-Thomason condition is stronger than $D_{\operatorname{qc}}$ being singly compactly generated.
1.2. Why it matters
Before stating our results, we briefly explain the importance of the existence of compact generators for $D_{\operatorname{qc}}$. Philosophically, $D_{\operatorname{qc}}$ and its subcategories encode homological information that is expected to reflect the geometry of the underlying algebraic stack. In particular, a substantial body of work has been developed around compact objects (see e.g. [Nee92, Nee96]). This provides a manageable framework for extracting and organizing such information. These ideas have played a central role in resolving conjectures [HR17b, Nee21, Nee24] and in the study of singularity theory for schemes (see e.g. [LV26]). They have also been essential in classification results for triangulated and derived categories (see e.g. [Hal16, Lan25]) and in $K$-theoretic applications (see e.g. [Jos02a, Jos02b, Toe99, Kri09]).
1.3. What is done
The main result of this paper is the following.
Theorem 1.1 (see Corollary˜4.7).
Let $\mathcal{X}$ be a concentrated regular algebraic stack with quasi-finite and locally separated diagonal. Then $\mathcal{X}$ satisfies the $1$-Thomason condition.
Particularly, Theorem˜1.1 asserts that $D_{\operatorname{qc}}(\mathcal{X})$ is singly compactly generated, and every closed subset of $|\mathcal{X}|$ arises as the support of a perfect complex. Moreover, Theorem˜1.1 improves [Hal22, Theorem 2.1] to single compact generation in the special case of quasi-finite and locally separated diagonal. Furthermore, we prove a more general statement where $\mathcal{X}$ does not need to be concentrated; see Theorem˜4.6. Here, ‘concentrated’ is a mild condition ensuring that compact objects coincide with perfect complexes. In our context, concentratedness coincides with being ‘tame’ in the sense of [Hal16]; see e.g. [DLM25, Proposition A.1].
In contrast to [HR17a], our proof avoids étale dévissage. Instead, the proof of Theorem˜1.1 proceeds in two steps. First, using finite flat presentations of Rydh [Ryd11], we reduce to descent along an étale morphism that is finite over a closed substack. Second, we apply recollement arguments for triangulated categories to glue compact generators; see Proposition˜3.1 for a construction.
Our result extends [HR17a, Theorem A] from quasi-finite separated diagonal to quasi-finite locally separated diagonal under regularity and concentratedness hypotheses. This yields genuinely new cases even among tame Deligne–Mumford stacks, notably allowing non-separated diagonal and arbitrary characteristic.
As a consequence, for such algebraic stacks $\mathcal{X}$ in Theorem˜1.1, the natural functor $D(\operatorname{Qcoh}(\mathcal{X}))\to D_{\operatorname{qc}}(\mathcal{X})$ is an equivalence (see [HNR19, Theorem 1.2]). Furthermore, for such algebraic stacks $\mathcal{X}$ in Theorem˜1.1, the underlying topological space $|\mathcal{X}|$ determines that of the associated Balmer spectra of $D_{\operatorname{qc}}(\mathcal{X})$ (in the sense of [Bal05]); see [HR17b, Theorem C].
On a separate note, we show that the $\beta$-Thomason condition descends along a finite, flat, surjective, morphism between Noetherian algebraic stacks. See Proposition˜A.1. We record the statement mainly for its independent interest and the simplicity of the argument. In fact, [HR17a, Theorem 6.6] proves a similar statement for the case of $\beta$-crispness.
1.4. What is next
Hall and Rydh introduced the notion of $\beta$-crispness in [HR17a]. This notion requires that the Thomason $\beta$-Thomason condition be stable ascending under étale, separated, quasi-compact morphisms representable by algebraic spaces. It is not clear whether $\mathcal{S}$ in Theorem˜1.1 is $1$-crisp. In particular, the inductive argument used in the proof of Theorem˜1.1 does not appear to interact well with immersions in the sense required for $\beta$-crispness. To be precise, any étale cover of $\mathcal{S}$ need not arise as an étale cover of $\mathcal{X}$.
Acknowledgments.
Lank was supported by ERC Advanced Grant 101095900-TriCatApp. The author greatly appreciates discussions with Timothy De Deyn, Jack Hall, Michal Hrbek, Kabeer Manali-Rahul, Fei Peng, and David Rydh.
2. Algebraic stacks
Our conventions for algebraic stacks are those of [Sta25]. For the derived pullback and pushforward adjunction, we adopt the conventions of [HR17a, §1] and [Ols07, LO08a, LO08b]. Unless otherwise specified, symbols such as $X$, $Y$, etc. denote schemes or algebraic spaces, while $\mathcal{X}$, $\mathcal{Y}$, etc. denote algebraic stacks. In this section, let $\mathcal{X}$ be a quasi-compact and quasi-separated algebraic stack.
Categories
We specify the triangulated categories that appear in our work. Let $\operatorname{Mod}(\mathcal{X})$ denote the Grothendieck abelian category of sheaves of $\mathcal{O}_{\mathcal{X}}$-modules on the lisse-étale site of $\mathcal{X}$. Define $\operatorname{Qcoh}(\mathcal{X})$ to be the strictly full subcategory (i.e. full and closed under isomorphisms) of $\operatorname{Mod}(\mathcal{X})$ consisting of quasi-coherent sheaves. Set $D(\mathcal{X}):=D(\operatorname{Mod}(\mathcal{X}))$ for the derived category of $\operatorname{Mod}(\mathcal{X})$. Denote by $D_{\operatorname{qc}}(\mathcal{X})$ the full subcategory of $D(\mathcal{X})$ consisting of complexes with quasi-coherent cohomology sheaves. Finally, let $\operatorname{Perf}(\mathcal{X})$ denote the full subcategory of perfect complexes in $D_{\operatorname{qc}}(\mathcal{X})$.
Support
Let $M\in\operatorname{Qcoh}(\mathcal{X})$. Set $\operatorname{supp}(M):=p\big(\operatorname{supp}(p^{\ast}M)\big)\subseteq|\mathcal{X}|$ where $p\colon U\to\mathcal{X}$ is any smooth surjective morphism from a scheme. One checks that this definition is independent of the choice of $p$. Now, given any $E\in D_{\operatorname{qc}}(\mathcal{X})$, let
$\operatorname{supp}(E):=\bigcup_{j\in\mathbb{Z}}\operatorname{supp}\big(\mathcal{H}^{j}(E)\big)\subseteq|\mathcal{X}|.$
This subset of $|\mathcal{X}|$ is called the support of $E$.
Concentratedness
A quasi-compact quasi-separated morphism of algebraic stacks is called concentrated if for every base change along a quasi-compact quasi-separated morphism, the derived pushforward has finite cohomological dimension. For instance, by [HR17a, Lemma 2.5(3)], morphisms which are representable by algebraic spaces are concentrated. An algebraic stack is concentrated if it is quasi-compact quasi-separated, and its structure morphism to $\operatorname{Spec}(\mathbb{Z})$ is concentrated. In fact, a quasi-compact quasi-separated algebraic stack $\mathcal{X}$ is concentrated if, and only if, any (hence all) of the following equivalent conditions hold: $\mathbf{R}\Gamma\colon D_{\operatorname{qc}}(\mathcal{X})\to D_{\operatorname{qc}}(\operatorname{Spec}(\mathbb{Z}))$ commutes with small coproducts; $\operatorname{Perf}(\mathcal{X})=D_{\operatorname{qc}}(\mathcal{X})^{c}$; or $\mathcal{O}_{\mathcal{X}}\in D_{\operatorname{qc}}(\mathcal{X})^{c}$. See [HR17a, §2, Lemma 2.5(5), & Remark 4.6] for details.
Perfect complexes
Perfect complexes may be defined on any ringed site [Sta25, Tag 08G4], in particular on the lisse-étale site of $\mathcal{X}$. A complex is strictly perfect if it is a bounded complex whose terms are direct summands of finite free modules; it is perfect if it is locally strictly perfect. Let $\operatorname{Perf}(\mathcal{X})$ denote the triangulated subcategory of $D_{\operatorname{qc}}(\mathcal{X})$ consisting of perfect complexes. In general, the compact objects of $D_{\operatorname{qc}}(\mathcal{X})$ are perfect complexes [HR17a, Lemma 4.4], although the converse need not hold. The two notions coincide precisely when the algebraic stack $\mathcal{X}$ is concentrated. Any compact object of $D_{\operatorname{qc}}(\mathcal{X})$ is a perfect complex and the support of a perfect complex has quasi-compact complement (see [HR17a, Lemmas 4.4 & 4.8]).
Thomason condition
In general, $D_{\operatorname{qc}}(\mathcal{X})$ need not be compactly generated (for instance, this fails for $D_{\operatorname{qc}}(B_{k}\mathbb{G}_{a})$ when $k$ is a field of positive characteristic; see [HNR19, Proposition 3.1]). A related notion is the ‘Thomason condition’ which was introduced in [HR17a]. We say that $\mathcal{X}$ satisfies the $\beta$-Thomason condition, for some cardinal $\beta$, if $D_{\operatorname{qc}}(\mathcal{X})$ is compactly generated by a set of cardinality at most $\beta$, and if for every closed subset $Z\subseteq|\mathcal{X}|$ with quasi-compact complement there exists a perfect complex $P\in\operatorname{Perf}(\mathcal{X})$ such that $\operatorname{supp}(P)=Z$. For example, any quasi-compact quasi-separated algebraic spaces satisfies the $1$-Thomason condition (see e.g. [Sta25, Tag 08HP]).
Approximation by compacts
Let $(T,E,m)$ be a triple consisting of a closed subset $T\subseteq|\mathcal{X}|$, $E\in D_{\operatorname{qc}}(\mathcal{X})$, and $m\in\mathbb{Z}$. We say that approximation by compacts holds for $(T,E,m)$ if there exists a $P\in D_{\operatorname{qc}}(\mathcal{X})^{c}$ with support contained in $T$ and a morphism $P\to E$ such $\mathcal{H}^{i}(P)\to\mathcal{H}^{i}(E)$ is an isomorphism for $i>m$ and surjective when $i=m$. More generally, we say $\mathcal{X}$ satisfies approximation by compacts if for every closed subset $T\subseteq|\mathcal{X}|$ with $\mathcal{X}\setminus T\to\mathcal{X}$ quasi-compact there exists an integer $r$ such that for any triple $(T,E,m)$ with $E$ is (mr)(m-r)-pseudocoherent (see e.g. [Sta25, Tag 08FT]) and $\mathcal{H}^{i}(E)$ is supported on $T$ for $i\geq m-r$, approximation by compacts holds. See [HLLP25, $\S 3$] for details. For example, any quasi-compact quasi-separated algebraic stack with quasi-finite and separated diagonal satisfies approximation by compacts [HLLP25, Corollary 5.4].
3. Recollements
We briefly recall the notion of a recollement. See [BBDG18, §1.4] for details. A recollement is a commutative diagram of triangulated categories and exact functors of the form
(3.1) $\mathcal{T}$$\mathcal{K}$$\mathcal{D}$$\scriptstyle{I}$$\scriptstyle{I_{\lambda}}$$\scriptstyle{I_{\rho}}$$\scriptstyle{Q}$$\scriptstyle{Q_{\lambda}}$$\scriptstyle{Q_{\rho}}$
satisfying:
  • $I_{\lambda}\dashv I\dashv I_{\rho}$ and $Q_{\lambda}\dashv Q\dashv Q_{\rho}$ (i.e. adjoint triples)
  • $I,Q_{\lambda},Q_{\rho}$ are fully faithful
  • $\ker(Q)$ coincides with the strictly full subcategory on objects of the form $I(T)$ where $T\in\mathcal{T}$.
In such a case, there are distinguished triangles
$(Q_{\lambda}\circ Q)(E)$ $\to E\to(I\circ I_{\lambda})(E)\to(Q_{\lambda}\circ Q)(E)[1],$
$(I\circ I_{\rho})(E)\to E\to(Q_{\rho}\circ Q)(E)\to(I\circ I_{\rho})(E)[1]$
which are functorial in $\mathcal{K}$. In particular, the natural transformations between these functors are given by the (co)units of the relevant adjoint pairs. Since $_{\lambda}, QQ$, $I$, and $I_{\lambda}$ are left adjoints, they preserve coproducts.
Proposition 3.1.
Let $\mathcal{X}$ be a quasi-compact quasi-separated algebraic stack. Suppose $j\colon\mathcal{U}\to\mathcal{X}$ is a quasi-compact open immersion. Set $Z:=|\mathcal{X}|\setminus|\mathcal{U}|$. There exists a recollement
$D_{\operatorname{qc}}(\mathcal{U})$$D_{\operatorname{qc}}(\mathcal{X})$$D_{\operatorname{qc},Z}(\mathcal{X})$$\scriptstyle{\mathbf{R}j_{\ast}}$$\scriptstyle{\mathbf{L}j^{\ast}}$$\scriptstyle{j^{!}}$$\scriptstyle{i^{!}}$$\scriptstyle{i_{\ast}}$$\scriptstyle{i^{\ast}}$
where in particular $i_{\ast}$ is the natural inclusion and $j^{!}$ the right adjoint of $\mathbf{R}j_{\ast}$.
Proof.
First, we spell out the existence of the needed functors. By [HR17a, Theorem 4.14(1)], the right adjoint of $\mathbf{R}j_{\ast}$ exists, which we denote by $j^{!}$. Now, there is a Verdier localization
$D_{\operatorname{qc},Z}(\mathcal{X})\xrightarrow{i_{\ast}}D_{\operatorname{qc}}(\mathcal{X})\xrightarrow{\mathbf{L}j^{\ast}}D_{\operatorname{qc}}(\mathcal{U}).$
This can be seen using e.g. [HR17b, Example 1.2]. Note that $\mathbf{R}j_{\ast}$ is right adjoint to $\mathbf{L}j^{\ast}$ on $D_{\operatorname{qc}}$ (see e.g. [HR17a, §1.3]). So, by [GP18, Lemma 2.2(ii)], $i_{\ast}$ must admit a right adjoint as well, which we denote by $i^{!}$ (loc. cit. pulls from [CPS88b, Theorem 1.1] and [CPS88a, Theorem 2.1]). It follows that
$D_{\operatorname{qc},Z}(\mathcal{X})\xleftarrow{i^{!}}D_{\operatorname{qc}}(\mathcal{X})\xleftarrow{\mathbf{R}j_{\ast}}D_{\operatorname{qc}}(\mathcal{U})$
is a Verdier localization sequence. However, being that $j^{!}$ is the right adjoint of $\mathbf{R}j_{\ast}$ on $D_{\operatorname{qc}}$, we can apply [GP18, Lemma 2.2(ii)] once more to see that $i^{!}$ admits a right adjoint as well, which we denote by $i^{\ast}$. Tying things together, we have the required data for a recollement; i.e. a Verdier localization sequence which is a localization and colocalization sequence in the sense of [Kra10, §4]. ∎
Remark 3.2.
Consider the recollement in Proposition˜3.1. In this case, $i^{!}$ preserves small coproducts, and so, $i_{\ast}$ preserves compact objects (see e.g. the proof of $\implies$ in [Nee96, Theorem 5.1]; which this fact does not require compact generation). Then, from [HR17a, Lemma 4.4(1)], it follows that any compact object of $D_{\operatorname{qc},Z}(\mathcal{X})$ must belong to $\operatorname{Perf}(\mathcal{X})$.
4. Proofs
This section proves our results. To start, we need a few lemmas.
Lemma 4.1.
Let $f\colon\mathcal{Y}\to\mathcal{X}$ be a morphism of quasi-compact quasi-separated algebraic stacks. Consider a closed subset $Z\subseteq|\mathcal{X}|$. Then
$\mathbf{L}f^{\ast}(D_{\operatorname{qc},Z}(\mathcal{X}))\subseteq D_{\operatorname{qc},f^{-1}(Z)}(\mathcal{Y}).$
Proof.
Let $p\colon U\to\mathcal{X}$ be a smooth surjective morphism from an affine scheme. Denote by $p^{\prime}\colon\mathcal{Y}\times_{\mathcal{X}}U\to\mathcal{Y}$ for the base change of $p$ along $f$. Choose a smooth surjective morphism $t\colon V\to\mathcal{Y}\times_{\mathcal{X}}U$ from an affine scheme. Note that $p^{\prime}\circ t$ is a smooth surjective morphism of finite presentation from a scheme. So, the desired claim follows if we check that
$\mathbf{L}(f\circ p^{\prime}\circ t)^{\ast}(D_{\operatorname{qc},Z}(\mathcal{X}))\subseteq D_{\operatorname{qc},(f\circ p^{\prime}\circ t)^{-1}(Z)}(V).$
Hence, we can reduce to the case of affine schemes. In this setting, $D_{\operatorname{qc},Z}(\mathcal{X})$ is compactly generated (see e.g. [Rou08, Theorem 6.8]). From Remark˜3.2, we know that $D_{\operatorname{qc},Z}(\mathcal{X})$ is compactly generated by a subset $\mathcal{B}$ of $D_{\operatorname{qc},Z}(\mathcal{X})\cap\operatorname{Perf}(\mathcal{X})$. Furthermore, every object of $D_{\operatorname{qc},Z}(\mathcal{X})$ is a homotopy colimit of iterated extensions of small coproducts of shifts of objects in $\mathcal{B}$ (see [Sta25, Tag 09SN]). So, the desired claim can be checked if we can show $\mathbf{L}f^{\ast}\mathcal{B}\subseteq D_{\operatorname{qc},f^{-1}(Z)}(\mathcal{Y})$. However, [HR17a, Lemma 4.8(2)] tells us
$\operatorname{supp}(\mathbf{L}f^{\ast}P)=f^{-1}(\operatorname{supp}(P))\subseteq f^{-1}(Z),$
which completes the proof. ∎
Lemma 4.2.
Let $F\colon\mathcal{T}\leftrightarrows\mathcal{S}\colon G$ be an adjoint pair of exact functors between triangulated categories admitting small coproducts. Assume $\mathcal{T}$ is compactly generated by a collection $\mathcal{B}$ and $G$ commutes with small coproducts. Then $G$ is conservative (i.e. $G(A)\cong 0\implies A\cong 0$) if, and only if, $F(\mathcal{B})$ compactly generates $\mathcal{S}$. In such a case, if $\mathcal{T}$ is compactly generated by a set of cardinality $\leq\beta$ for some cardinal $\beta$, then $\mathcal{B}$ satisfies the same condition.
Proof.
This is known to some but we spell it out for convenience. That $G$ commutes with small coproducts ensures that $F(\mathcal{T}^{c})\subseteq\mathcal{S}^{c}$ (see e.g. the proof of $\implies$ in [Nee96, Theorem 5.1]; which this fact does not require $\mathcal{T}$ to be compactly generated). First, assume $G$ is conservative. Let $E\in\mathcal{B}$ satisfy $\operatorname{Hom}(F(B),E[n])\cong 0$ for all $B\in\mathcal{B}$ and $n\in\mathbb{Z}$. From adjunction, it follows that $\operatorname{Hom}(B,G(E)[n])\cong 0$. As $\mathcal{B}$ compactly generates $\mathcal{T}$, it follows that $G(E)\cong 0$. However, $G$ being conservative implies $E\cong 0$. So, $F(\mathcal{B})$ compactly generates $\mathcal{S}$.
Conversely, assume that $F(\mathcal{B})$ compactly generates $\mathcal{S}$. Let $E\in\mathcal{S}$ such that $G(E)\cong 0$. By adjunction, it follows that $0\cong\operatorname{Hom}(F(B),E[n])$ for all $B\in\mathcal{B}$ and $n\in\mathbb{Z}$. Yet, the assumption implies $E\cong 0$. Hence, $G$ must be conservative.
That the last claim holds follows from the proof above. ∎
Lemma 4.3.
Let $t\colon\mathcal{Y}\to\mathcal{X}$ be an immersion of quasi-compact quasi-separated algebraic stacks. If $\mathcal{X}$ satisfies the $\beta$-Thomason condition, then so does $\mathcal{Y}$.
Proof.
This is [Nee23, Observation 5.6], but we give a slightly different proof. An immersion is quasi-affine, and hence, concentrated. So, [HR17a, Lemma 8.2] guarantees that $D_{\operatorname{qc}}(\mathcal{Y})$ is compactly generated by a collection of at most $\leq\beta$ objects. Here, $t$ is injective on the underlying topological spaces. Let $Z$ be a closed subset of $|\mathcal{Y}|$. Set $Z^{\prime}$ to be the closure of $t(Z)$ in $|\mathcal{X}|$. Then $t^{-1}(Z^{\prime})=Z$. Moreover, the hypothesis on $\mathcal{X}$ allows us to find a $P\in\operatorname{Perf}(\mathcal{X})$ such that $\operatorname{supp}(P)=Z^{\prime}$. However, [HR17a, Lemma 4.8(2)] tells us $t^{-1}(\operatorname{supp}(P))=\operatorname{supp}(\mathbf{L}t^{\ast}P)$, which completes the proof because $\mathbf{L}t^{\ast}P\in\operatorname{Perf}(\mathcal{Y})$. ∎
Lemma 4.4.
Let $f\colon\mathcal{Y}\to\mathcal{X}$ be a finite morphism of Noetherian algebraic stacks. If $D_{\operatorname{qc}}(\mathcal{Y})$ is compactly generated by $\mathcal{B}\subseteq D_{\operatorname{qc}}(\mathcal{Y})^{c}$, then the adjoint pair $\mathbf{R}f_{\ast}$ and $f^{!}$ restricts to $D_{\operatorname{qc}}(\mathcal{Y})$ and $D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X})$. Additionally, if $\mathbf{R}f_{\ast}\mathcal{O}_{\mathcal{Y}}\in\operatorname{Perf}(\mathcal{X})$, then $\mathbf{R}f_{\ast}\mathcal{B}$ compactly generates $D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X})$.
Proof.
To start, we show the claim regarding restricting the adjunction. Recall, by [HR17a, Theorem 4.14(1)], the right adjoint of $\mathbf{R}f_{\ast}$ on $D_{\operatorname{qc}}$ exists because $f$ is concentrated. It is easy to see that $f^{!}(D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X}))\subseteq D_{\operatorname{qc}}(\mathcal{Y})$. So, we only need to show that $\mathbf{R}f_{\ast}(D_{\operatorname{qc}}(\mathcal{Y}))\subseteq D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X})$. Let $E\in D_{\operatorname{qc}}(\mathcal{Y})$. Since $D_{\operatorname{qc}}(\mathcal{Y})$ is compactly generated by $\mathcal{B}$, [Kra22, Corollary 3.4.9] (or [Sta25, Tag 09SN]) tells us $E$ can be written as the homotopy colimit of a sequence
$E_{0}\xrightarrow{\phi_{0}}E_{1}\xrightarrow{\phi_{1}}E_{2}\xrightarrow{\phi_{2}}\cdots$
such that $E_{0}\cong 0$ and $\operatorname{cone}(\phi_{i})$ is a small coproduct of shifts of objects in $\mathcal{B}$. Now, as $\mathbf{R}f_{\ast}$ preserves homotopy colimits, it suffices to show $\mathbf{R}f_{\ast}E_{i}$ and $\operatorname{cone}(\mathbf{R}f_{\ast}\phi_{i})$ belong to $D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X})$. Using an inductive argument on $i$, it suffices to show small coproducts of shifts of objects from $\mathcal{B}$ satisfy the same condition. Furthermore, $\mathbf{R}f_{\ast}$ preserving small coproducts and shifts, so we can reduce to checking the same condition on just the objects of $\mathcal{B}$. However, this can be seen by reducing to schemes (see e.g. [GW23, Remark 23.46(2)]), e.g. $\mathbf{R}f_{\ast}\mathcal{B}\subseteq D^{b}_{\operatorname{coh},f(|\mathcal{Y}|)}(\mathcal{X})$.
Lastly, we show that $\mathbf{R}f_{\ast}\mathcal{B}$ compactly generates $D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X})$. Since [HR17a, Theorem 4.14(4)] tells us $f^{!}$ preserves small coproducts, Lemma˜4.2 ensures that it suffices to show that $f^{!}\colon D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X})\to D_{\operatorname{qc}}(\mathcal{Y})$ is conservative. So, let $E\in D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X})$ satisfy $f^{!}E\cong 0$. By [HR17a, Theorem 4.14(2)], it follows that
$0\cong\mathbf{R}f_{\ast}f^{!}E\cong\operatorname{\mathbf{R}\mathcal{H}\!\mathit{om}}(\mathbf{R}f_{\ast}\mathcal{O}_{\mathcal{Y}},E).$
However, $\mathbf{R}f_{\ast}\mathcal{O}_{\mathcal{Y}}\in\operatorname{Perf}(\mathcal{X})\cap D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X})$, and so, [HR17a, Lemma 4.9] ensures that $E\cong 0$. Hence, we see that the restriction of $f^{!}$ on $D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X})$ is conservative. ∎
Remark 4.5.
Recall that a collection $\mathcal{B}$ in a category $\mathcal{C}$ is said to generate $\mathcal{C}$ if for any nonzero $E\in\mathcal{C}$ there is an $n\in\mathbb{Z}$ and $B\in\mathcal{B}$ such that $\operatorname{Hom}(B[n],E)\not\cong 0$.
Theorem 4.6.
Let $\mathcal{X}$ be a quasi-compact quasi-separated regular algebraic stack with quasi-finite and locally separated diagonal. Then there is an object of $\operatorname{Perf}(\mathcal{X})$ which generates $D_{\operatorname{qc}}(\mathcal{X})$.
Proof.
Using [Ryd11, Theorem 7.2], we can find an étale, surjective, finitely presented (i.e. of finite presentation), representable by algebraic spaces morphism $p\colon\mathcal{Y}\to\mathcal{X}$ and a finite, flat, surjective, finitely presented morphism $v\colon U\to\mathcal{Y}$ from a quasi-affine scheme. Furthermore, appealing to [Ryd11, Proposition 4.4], there is a finite sequence of quasi-compact open immersions
$\emptyset=:\mathcal{X}_{0}\xrightarrow{j_{0}}\mathcal{X}_{1}\xrightarrow{j_{1}}\cdots\xrightarrow{j_{n-2}}\mathcal{X}_{n-1}\xrightarrow{j_{n-1}}\mathcal{X}_{n}=:\mathcal{X}$
such that the base change of $f$ to $|\mathcal{X}_{c}|\setminus|\mathcal{X}_{c-1}|$ (when given any closed substack structure) is étale, finite, and of constant rank. Note that every algebraic stack above is Noetherian.
By inducting on $n$, we can reduce to the setting where $f$ is finite after base change along a closed immersion $t\colon\mathcal{Z}\to\mathcal{X}$ and the open substack $\mathcal{U}:=|\mathcal{X}|\setminus t(|\mathcal{Z}|)$ satisfies $D_{\operatorname{qc}}(\mathcal{U})$ being generated by an object $G_{U}\in\operatorname{Perf}(\mathcal{U})$. Set $Z=t(|\mathcal{Z}|)$ and let $j\colon\mathcal{U}\to\mathcal{X}$ be the associated open immersion. Now, consider the commutative diagram obtained by base changes,
[…]
where each internal square is fibered. Here, $f^{\prime}$ and $v^{\prime}$ are flat, finite, and surjective morphisms. Since $t^{\prime\prime}$ is closed immersion, we have that $U\times_{\mathcal{X}}\mathcal{Z}$ is a Noetherian scheme. So, there is a $P\in\operatorname{Perf}(U\times_{\mathcal{X}}\mathcal{Z})$ which compactly generates $D_{\operatorname{qc}}(U\times_{\mathcal{X}}\mathcal{Z})$. Then Lemma˜4.4 (or Proposition˜A.1) tells us that $\mathbf{R}(f^{\prime}\circ v^{\prime})_{\ast}P$ compactly generates $D_{\operatorname{qc}}(\mathcal{Z})$ as $f^{\prime}\circ v^{\prime}$ is flat, finite, and surjective.
By [HLLP25, Proposition B.1], there is a Verdier localization sequence
$D^{b}_{\operatorname{coh},Z}(\mathcal{X})\xrightarrow{i_{\ast}}D^{b}_{\operatorname{coh}}(\mathcal{X})\xrightarrow{\mathbf{L}j^{\ast}}D^{b}_{\operatorname{coh}}(\mathcal{U}).$
However, [DLMP25, Theorem 3.7] tells us that $\operatorname{Perf}=D^{b}_{\operatorname{coh}}$ in each case. This allows us to find a $G\in\operatorname{Perf}(\mathcal{X})$ such that $\mathbf{L}j^{\ast}G\cong G_{U}$; that is, $\mathbf{L}j^{\ast}G$ generates $D_{\operatorname{qc}}(\mathcal{U})$ and belongs to $\operatorname{Perf}(\mathcal{U})$. We claim that $G\oplus\mathbf{R}(f^{\prime}\circ v^{\prime})_{\ast}P$ generates $D_{\operatorname{qc}}(\mathcal{X})$. Towards this end, consider the recollement obtained in Proposition˜3.1,
$D_{\operatorname{qc}}(\mathcal{U})$$D_{\operatorname{qc}}(\mathcal{X})$$D_{\operatorname{qc},Z}(\mathcal{X})$$\scriptstyle{\mathbf{R}j_{\ast}}$$\scriptstyle{\mathbf{L}j^{\ast}}$$\scriptstyle{j^{!}}$$\scriptstyle{i^{!}}$$\scriptstyle{i_{\ast}}$$\scriptstyle{i^{\ast}}$
where in particular $i_{\ast}$ is the natural inclusion and $j^{!}$ the right adjoint of $\mathbf{R}j_{\ast}$. Let $E\in D_{\operatorname{qc}}(\mathcal{X})$ satisfy $\operatorname{Hom}(G\oplus\mathbf{R}(f^{\prime}\circ v^{\prime})_{\ast}P[n],E)=0$ for all $n\in\mathbb{Z}$. Since $i_{\ast}$ is the inclusion, we have symbolically that $\mathbf{R}(f^{\prime}\circ v^{\prime})_{\ast}P=i_{\ast}\mathbf{R}(f^{\prime}\circ v^{\prime})_{\ast}P$. From the recollement above, there is a distinguished triangle
$i_{\ast}i^{!}E\to E\to\mathbf{R}j_{\ast}\mathbf{L}j^{\ast}E\to i_{\ast}i^{!}E[1].$
Here, in such a case, we have that $\operatorname{Hom}(i_{\ast}\mathbf{R}(f^{\prime}\circ v^{\prime})_{\ast}P[n],E)\cong 0$ for all $n\in\mathbb{Z}$. So, adjunction tells us $\operatorname{Hom}(\mathbf{R}(f^{\prime}\circ v^{\prime})_{\ast}P[n],i^{!}E)\cong 0$ for all $n\in\mathbb{Z}$. However, $\mathbf{R}(f^{\prime}\circ v^{\prime})_{\ast}P$ generates $D_{\operatorname{qc},Z}(\mathcal{X})$, so $i^{!}E\cong 0$. This tells us the morphism $E\to\mathbf{R}j_{\ast}\mathbf{L}j^{\ast}E$ from the distinguished triangle above is an isomorphism. Once more, from adjunction, we then have that $\operatorname{Hom}(G[n],\mathbf{R}j_{\ast}\mathbf{L}j^{\ast}E)=0$ for all $n\in\mathbb{Z}$. So, $\operatorname{Hom}(\mathbf{L}j^{\ast}G[n],\mathbf{L}j^{\ast}E)=0$ via adjunction for any such $n$. However, $\mathbf{L}j^{\ast}G$ generates $D_{\operatorname{qc}}(\mathcal{U})$, so $\mathbf{L}j^{\ast}E\cong 0$. Consequently, $E\cong 0$, which completes the proof. ∎
Corollary 4.7.
Let $\mathcal{X}$ be a concentrated regular algebraic stack with quasi-finite and locally separated diagonal. Then $\mathcal{X}$ satisfies the $1$-Thomason condition. Moreover, if $\mathcal{S}\to\mathcal{X}$ is an immersion, then $\mathcal{S}$ satisfies the $1$-Thomason condition.
Proof.
That the last claim holds if we can prove the first by Lemma˜4.3. As $\mathcal{X}$ is regular and Noetherian, it follows that $\operatorname{Perf}(\mathcal{X})=D^{b}_{\operatorname{coh}}(\mathcal{X})$ (see e.g. [DLMP25, Theorem 3.7]). Let $W\subseteq|\mathcal{X}|$ be closed. Denote by $i\colon\mathcal{W}\to\mathcal{X}$ for the closed immersion from the reduced induced closed substack structure on $W$. If $\mathcal{X}$ is concentrated, then $\mathbf{R}i_{\ast}\mathcal{O}_{\mathcal{W}}\in D_{\operatorname{qc}}(\mathcal{X})^{c}$ and has support coinciding with $W$. Moreover, by Theorem˜4.6, there is an $E\in D^{b}_{\operatorname{coh}}(\mathcal{X})$ which generates $D_{\operatorname{qc}}(\mathcal{X})$. If $\mathcal{X}$ is concentrated, then $\operatorname{Perf}(\mathcal{X})=D_{\operatorname{qc}}(\mathcal{X})^{c}$, and so $D_{\operatorname{qc}}(\mathcal{X})$ is compactly generated as desired. ∎
Appendix A Descending Thomason condition
While not needed in proving our main results, we record the following for sake of interest (cf. [HR17a, Theorem 6.6] in the case for $\beta$-crispness).
Proposition A.1.
Let $f\colon\mathcal{Y}\to\mathcal{X}$ be a finite, flat, surjective, morphism between Noetherian algebraic stacks. If $\mathcal{Y}$ satisfies the $\beta$-Thomason condition, then so does $\mathcal{X}$.
Proof.
Since $f\colon\mathcal{Y}\to\mathcal{X}$ is a finite, flat, surjective morphism of Noetherian algebraic stacks, it follows that $\mathbf{R}f_{\ast}\mathcal{O}_{\mathcal{Y}}\in\operatorname{Perf}(\mathcal{X})$. Denote by $f^{!}$ the right adjoint of $\mathbf{R}f_{\ast}$ on $D_{\operatorname{qc}}$. In our case, [HR17a, Theorem 4.14] tells us that $f^{!}$ preserves small coproducts and is conservative. By the hypothesis, we know that there exists a collection $\mathcal{B}$ (of some cardinality $\leq\beta$) which compactly generates $D_{\operatorname{qc}}(\mathcal{Y})$.
Choose a closed $Z\subseteq|\mathcal{X}|$. The hypothesis on $\mathcal{Y}$ tells us there is a $P\in\operatorname{Perf}(\mathcal{Y})$ such that $\operatorname{supp}(P)=f^{-1}(Z)$. So, appealing to [HR17a, Lemma 4.10(2)], we know that $D_{\operatorname{qc},f^{-1}(Z)}(\mathcal{Y})$ is compactly generated by objects of the form $B\otimes^{\mathbf{L}}P$ with $B\in\mathcal{B}$.
We claim that the collection of $\mathbf{R}f_{\ast}(B\otimes^{\mathbf{L}}P)$ with $B\in\mathcal{B}$ compactly generates $D_{\operatorname{qc},Z}(\mathcal{X})$. Towards that end, we show that the adjunction $\mathbf{R}f_{\ast}\dashv f^{!}$ restricts to an adjoint pair between $D_{\operatorname{qc},f^{-1}(Z)}(\mathcal{Y})$ and $D_{\operatorname{qc},Z}(\mathcal{X})$. By [HR17a, Theorem 4.14(4)], we know that $f^{!}\mathcal{O}_{\mathcal{X}}\otimes^{\mathbf{L}}\mathbf{L}f^{\ast}E\cong f^{!}E$ for all $E\in D_{\operatorname{qc},Z}(\mathcal{X})$. This ensures that $\operatorname{supp}(f^{!}(E))\subseteq\operatorname{supp}(\mathbf{L}f^{\ast}E)$. Moreover, Lemma˜4.1 promises that $\operatorname{supp}(\mathbf{L}f^{\ast}E)\subseteq f^{-1}(Z)$, and so $f^{!}D_{\operatorname{qc},Z}(\mathcal{X})\subseteq D_{\operatorname{qc},f^{-1}(Z)}(\mathcal{Y})$.
Next, we check that $\mathbf{R}f_{\ast}D_{\operatorname{qc},f^{-1}(Z)}(\mathcal{Y})\subseteq D_{\operatorname{qc},Z}(\mathcal{X})$. However, being that $D_{\operatorname{qc}}(\mathcal{Z})$ is compactly generated by the collection $B\otimes^{\mathbf{L}}P$ with $B\in\mathcal{B}$, every object of $D_{\operatorname{qc}}(\mathcal{Z})$ is a homotopy colimit of iterated extensions of small coproducts of shifts of the collection $B\otimes^{\mathbf{L}}P$ (see [Sta25, Tag 09SN]). So, this can be checked by showing $\mathbf{R}f_{\ast}(B\otimes^{\mathbf{L}}P)\in D_{\operatorname{qc},Z}(\mathcal{X})$. However, this can be seen by reducing to schemes (see e.g. [GW23, Remark 23.46(2)]).
Now, we finish the proof. As we have shown the adjunction restricts, we can make use of the fact that $f^{!}$ is conservative by [HR17a, Theorem 4.14(4)]. Consequently, Lemma˜4.2 implies the collection of $\mathbf{R}f_{\ast}(B\otimes^{\mathbf{L}}P)$ with $B\in\mathcal{B}$ compactly generates $D_{\operatorname{qc},Z}(\mathcal{X})$. This tells us every object of $D_{\operatorname{qc},Z}(\mathcal{X})$ is a homotopy colimit of iterated extensions of small coproducts of shifts of the collection $\mathbf{R}f_{\ast}(B\otimes^{\mathbf{L}}P)$ with $B\in\mathcal{B}$ (see [Sta25, Tag 09SN]). So, if the support of every $\mathbf{R}f_{\ast}(B\otimes^{\mathbf{L}}P)$ is properly contained in $Z$, then so must the support of every object in $D_{\operatorname{qc},Z}(\mathcal{X})$, which is absurd. Therefore, there must exist some $B\in\mathcal{B}$ such that $\operatorname{supp}(\mathbf{R}f_{\ast}(B\otimes^{\mathbf{L}}P))=Z$. ∎