### v3

# Perfect generation for regular algebraic stacks

Abstract.
We show that the derived category of complexes with quasi-coherent cohomology on a regular Noetherian algebraic stack with quasi-finite diagonal is generated by a single perfect complex. In the concentrated case, the category is singly compactly generated. Key ingredients in the proofs include gluing generators along recollement and the use of suitable filtrations and presentations of the algebraic stack.
Key words and phrases:
Algebraic stacks, perfect complexes, Thomason condition, derived categories, compact objects
2020 Mathematics Subject Classification:
14A30 (primary), 14D23, 14F08, 18G80
1. Introduction
1.1. What is known
Consider the derived category $D_{\operatorname{qc}}$ of complexes with quasi-coherent cohomology on an algebraic stack. It is classical that $D_{\operatorname{qc}}$ is singly compactly generated for quasi-compact quasi-separated schemes [BV03, Theorem 3.1.1]. However, the situation is different for algebraic stacks, e.g. perfect complexes $P\in D_{\operatorname{qc}}$ need not be compact.
Early extensions for single object compact generation to various classes of algebraic stacks were obtained in [TOË12, Corollary 5.2], [KRI09, proof of Proposition 5.5], and [BFN10, §3.3]. More recently, the work of Hall–Rydh has pushed these results further. These include quasi-compact quasi-separated algebraic stacks with quasi-finite separated diagonal [HR17a, Theorem A] and quasi-compact quasi-separated Deligne–Mumford $\mathbb{Q}$-stacks [HR18, Theorem 7.4].
It was shown in [HAL22] that $D_{\operatorname{qc}}$ is compactly generated for any affine-pointed concentrated regular algebraic stack of finite Krull dimension. However, it is not clear whether single object compact generation occurs. Here, ‘concentrated’ is a mild condition ensuring that compact objects coincide with perfect complexes. In the case of finitely presented inertia, being concentrated wth quasi-finite diagonal coincides with being ‘tame’ in the sense of [HAL16]; see e.g. [DLM25b, Proposition A.1]. Also, ‘affine-pointed’ means that every morphism from the spectrum of a field is affine.
1.2. What is new
Our focus is on single object compact generation for regular algebraic stacks. As a first step, we consider a weaker notion of generation in triangulated categories. Recall that a collection $\mathcal{B}$ in a category $\mathcal{C}$ is said to generate $\mathcal{C}$ if for any nonzero $E\in\mathcal{C}$ there is an $n\in\mathbb{Z}$ and $B\in\mathcal{B}$ such that $\operatorname{Hom}(B[n],E)\not\cong 0$.
We establish generation by a single perfect complex.
Theorem 1.1.
Let $\mathcal{X}$ be a quasi-compact quasi-separated regular algebraic stack with quasi-finite diagonal. Then there is a $P\in\operatorname{Perf}(\mathcal{X})$ which generates $D_{\operatorname{qc}}(\mathcal{X})$.
Note that we do not assume separated diagonal or that $\mathcal{X}$ is concentrated. Moreover, Theorem˜1.1 yields genuinely new cases, even among Deligne–Mumford stacks in arbitrary characteristic under the regularity hypothesis (e.g. smooth over a field or DVR). It appears that generation by a single perfect complex which need not be compact seems quite new.
The proof proceeds in two steps. First, we use the presentations of [HR18] for algebraic stacks with finite stabilizers. This allows us to argue inductively on the length of a monomorphic splitting sequence of a Nisnevich covering. Second, we use recollement techniques for triangulated categories to glue generators (see Proposition˜3.1). We find the use of recollements here to be of independent interest.
This reduces the problem to detecting single object generation on suitable locally closed substacks. We then apply finite duality to show that subcategories of the form $D_{\operatorname{qc},Z}$ are generated by a single object, where the regularity hypothesis is used. In fact, the argument suggests single object compact generation, but we are not able to prove this (see Remark˜4.5).
A straightforward application is the following.
Corollary 1.2.
$D_{\operatorname{qc}}(\mathcal{X})$ is singly compactly generated for any concentrated regular algebraic stack $\mathcal{X}$ with quasi-finite diagonal.
Under the quasi-finite diagonal condition, this result improves [HAL22, Theorem 2.1] to single object compact generation and removes finiteness of Krull dimension. It applies to the case $\mathcal{X}$ is a regular Noetherian quasi-DM $\mathbb{Q}$-stack with affine stabilizers because concentratedness follows from [HR15, Theorem 2.1(1)]. Furthermore, it applies when $\mathcal{X}$ is smooth, finitely presented, and quasi-DM over a DVR (possibly of mixed characteristic) where the stabilizers of $\mathcal{X}$ are affine and nice (in the sense of [HR15]).
Acknowledgments.
Lank was supported by ERC Advanced Grant 101095900-TriCatApp. The author greatly appreciates comments and/or discussions with Jarod Alper, Timothy De Deyn, Jack Hall, Michal Hrbek, Kabeer Manali-Rahul, Fei Peng, and David Rydh.
2. Algebraic stacks
Our conventions for algebraic stacks are those of [STA26]. For the derived pullback and pushforward adjunction, we adopt the conventions of [HR17a, §1] and [OLS07, LO08a, LO08b]. Unless otherwise specified, symbols such as $X$, $Y$, etc. denote schemes or algebraic spaces, while $\mathcal{X}$, $\mathcal{Y}$, etc. denote algebraic stacks. In this section, let $\mathcal{X}$ be a quasi-compact quasi-separated algebraic stack.
Categories
We specify the triangulated categories that appear in our work. Let $\operatorname{Mod}(\mathcal{X})$ denote the Grothendieck abelian category of sheaves of $\mathcal{O}_{\mathcal{X}}$-modules on the lisse-étale site of $\mathcal{X}$. Define $\operatorname{Qcoh}(\mathcal{X})$ to be the strictly full subcategory (i.e. full and closed under isomorphisms) of $\operatorname{Mod}(\mathcal{X})$ consisting of quasi-coherent sheaves. Set $D(\mathcal{X}):=D(\operatorname{Mod}(\mathcal{X}))$ for the derived category of $\operatorname{Mod}(\mathcal{X})$. Denote by $D_{\operatorname{qc}}(\mathcal{X})$ the full subcategory of $D(\mathcal{X})$ consisting of complexes with quasi-coherent cohomology. Finally, let $\operatorname{Perf}(\mathcal{X})$ denote the full subcategory of perfect complexes in $D_{\operatorname{qc}}(\mathcal{X})$.
Support
Let $M\in\operatorname{Qcoh}(\mathcal{X})$. Set $\operatorname{supp}(M):=p\big(\operatorname{supp}(p^{\ast}M)\big)\subseteq|\mathcal{X}|$ where $p\colon U\to\mathcal{X}$ is any smooth surjective morphism from a scheme. One can check that this definition is independent of the choice of $p$. Now, given any $E\in D_{\operatorname{qc}}(\mathcal{X})$, let
$\operatorname{supp}(E):=\bigcup_{j\in\mathbb{Z}}\operatorname{supp}\big(\mathcal{H}^{j}(E)\big)\subseteq|\mathcal{X}|.$
This subset of $|\mathcal{X}|$ is called the support of $E$.
We record two lemmas for convenience.
Lemma 2.1.
Let $f\colon\mathcal{Y}\to\mathcal{X}$ be a morphism of quasi-compact quasi-separated algebraic stacks. Consider a closed $Z\subseteq|\mathcal{X}|$. Then
$\mathbf{L}f^{\ast}(D_{\operatorname{qc},Z}(\mathcal{X}))\subseteq D_{\operatorname{qc},f^{-1}(Z)}(\mathcal{Y}).$
Proof.
Let $p\colon U\to\mathcal{X}$ be a smooth surjective morphism from an affine scheme. Denote by $p^{\prime}\colon\mathcal{Y}\times_{\mathcal{X}}U\to\mathcal{Y}$ for the base change of $p$ along $f$. Choose a smooth surjective morphism $t\colon V\to\mathcal{Y}\times_{\mathcal{X}}U$ from an affine scheme. Note that $p^{\prime}\circ t$ is a smooth surjective morphism of finite presentation from a scheme. So, the desired claim follows if we check that
$\mathbf{L}(f\circ p^{\prime}\circ t)^{\ast}(D_{\operatorname{qc},Z}(\mathcal{X}))\subseteq D_{\operatorname{qc},(f\circ p^{\prime}\circ t)^{-1}(Z)}(V).$
Hence, we can reduce to the case of affine schemes. In this setting, $D_{\operatorname{qc},Z}(\mathcal{X})$ is compactly generated (see [ROU08, Theorem 6.8]). From Remark˜3.2, we know that $D_{\operatorname{qc},Z}(\mathcal{X})$ is compactly generated by a subset $\mathcal{B}$ of $D_{\operatorname{qc},Z}(\mathcal{X})\cap\operatorname{Perf}(\mathcal{X})$. Furthermore, every object of $D_{\operatorname{qc},Z}(\mathcal{X})$ is a homotopy colimit of iterated extensions of small coproducts of shifts of objects in $\mathcal{B}$ (see [STA26, Tag 09SN]). Therefore, the desired claim can be checked if we can show $\mathbf{L}f^{\ast}\mathcal{B}\subseteq D_{\operatorname{qc},f^{-1}(Z)}(\mathcal{Y})$. However, [HR17a, Lemma 4.8(2)] tells us
$\operatorname{supp}(\mathbf{L}f^{\ast}P)=f^{-1}(\operatorname{supp}(P))\subseteq f^{-1}(Z),$
which completes the proof. ∎
Lemma 2.2.
Let $f\colon\mathcal{Y}\to\mathcal{X}$ be a proper morphism of Noetherian algebraic stacks which is representable by scheme. For any closed subset $Z\subseteq|\mathcal{Y}|$, one has $\mathbf{R}f_{\ast}D_{\operatorname{qc},Z}(\mathcal{Y})\subseteq D_{\operatorname{qc},f(Z)}(\mathcal{X})$.
Proof.
Here, $f$ is concentrated. So, the problem is smooth local via [HR17a, Theorem 2.6(2)]. Hence, we can impose Noetherian schemes $X$ and $Y$. In such a case, $D_{\operatorname{qc},Z}(Y)$ is compactly generated by perfect complexes on $Y$ which have support contained in $Z$ (see [ROU08, Theorem 6.8]). Furthermore, every object of $D_{\operatorname{qc},Z}(Y)$ is a homotopy colimit of iterated extensions of small coproducts of shifts of objects in $\mathcal{B}$ (see [STA26, Tag 09SN]). Therefore, the desired claim can be checked if we can show $\mathbf{R}f_{\ast}\mathcal{B}\subseteq D_{\operatorname{qc},f(Z)}(X)$. However, this is known for proper morphisms of Noetherian schemes, see e.g. [GW23, Remark 23.46(2)]. ∎
Concentratedness
A quasi-compact quasi-separated morphism of algebraic stacks is called concentrated if for every base change along a quasi-compact quasi-separated morphism, the derived pushforward has finite cohomological dimension. For instance, by [HR17a, Lemma 2.5(3)], morphisms which are representable by algebraic spaces are concentrated. An algebraic stack is concentrated if it is quasi-compact quasi-separated, and its structure morphism to $\operatorname{Spec}(\mathbb{Z})$ is concentrated. In fact, a quasi-compact quasi-separated algebraic stack $\mathcal{X}$ is concentrated if, and only if, $\operatorname{Perf}(\mathcal{X})=D_{\operatorname{qc}}(\mathcal{X})^{c}$ if, and only if, $\mathcal{O}_{\mathcal{X}}\in D_{\operatorname{qc}}(\mathcal{X})^{c}$. See [HR17a, §2, Lemma 2.5(5), & Remark 4.6] for details.
Perfect complexes
Perfect complexes may be defined on any ringed site [STA26, Tag 08G4], in particular on the lisse-étale site of $\mathcal{X}$. A complex is strictly perfect if it is a bounded complex whose terms are direct summands of finite free modules; it is perfect if it is locally strictly perfect. Let $\operatorname{Perf}(\mathcal{X})$ denote the triangulated subcategory of $D_{\operatorname{qc}}(\mathcal{X})$ consisting of perfect complexes. In general, the compact objects of $D_{\operatorname{qc}}(\mathcal{X})$ are perfect complexes [HR17a, Lemma 4.4], although the converse need not hold. The two notions coincide precisely when the algebraic stack $\mathcal{X}$ is concentrated. Any compact object of $D_{\operatorname{qc}}(\mathcal{X})$ is a perfect complex and the support of a perfect complex has quasi-compact complement (see [HR17a, Lemmas 4.4 & 4.8]).
Thomason condition
In general, $D_{\operatorname{qc}}(\mathcal{X})$ need not be compactly generated (for instance, this fails for $D_{\operatorname{qc}}(B_{k}\mathbb{G}_{a})$ when $k$ is a field of positive characteristic; see [HNR19, Proposition 3.1]). A related notion is the ‘Thomason condition’ which was introduced in [HR17a]. We say that $\mathcal{X}$ satisfies the $\beta$-Thomason condition, for some cardinal $\beta$, if $D_{\operatorname{qc}}(\mathcal{X})$ is compactly generated by a set of cardinality at most $\beta$, and if for every closed subset $Z\subseteq|\mathcal{X}|$ with quasi-compact complement there exists a perfect complex $P\in\operatorname{Perf}(\mathcal{X})$ such that $\operatorname{supp}(P)=Z$. For example, any quasi-compact quasi-separated algebraic spaces satisfies the $1$-Thomason condition (see e.g. [STA26, Tag 08HP]).
3. Recollements
3.1. Reminder
We briefly recall the notion of a recollement. See [BBD+18, §1.4] for details. A recollement is a commutative diagram of triangulated categories and exact functors of the form
(3.1) $\mathcal{T}$$\mathcal{K}$$\mathcal{D}$$\scriptstyle{I}$$\scriptstyle{I_{\lambda}}$$\scriptstyle{I_{\rho}}$$\scriptstyle{Q}$$\scriptstyle{Q_{\lambda}}$$\scriptstyle{Q_{\rho}}$
satisfying:
  • $I_{\lambda}\dashv I\dashv I_{\rho}$ and $Q_{\lambda}\dashv Q\dashv Q_{\rho}$ (i.e. adjoint triples)
  • $I,Q_{\lambda},Q_{\rho}$ are fully faithful
  • $\ker(Q)$ coincides with the strictly full subcategory on objects of the form $I(T)$ where $T\in\mathcal{T}$.
In such a case, there are distinguished triangles
$(Q_{\lambda}\circ Q)(E)$ $\to E\to(I\circ I_{\lambda})(E)\to(Q_{\lambda}\circ Q)(E)[1],$
$(I\circ I_{\rho})(E)\to E\to(Q_{\rho}\circ Q)(E)\to(I\circ I_{\rho})(E)[1]$
which are functorial in $\mathcal{K}$. In particular, the natural transformations between these functors are given by the (co)units of the relevant adjoint pairs. Since $_{\lambda}, QQ$, $I$, and $I_{\lambda}$ are left adjoints, they preserve coproducts.
3.2. A special case
We record some results in setting of algebraic stacks. The following is well-known for schemes (see e.g. [JØR09, Theorem 1]).
Proposition 3.1.
Let $\mathcal{X}$ be a quasi-compact quasi-separated algebraic stack. Suppose $j\colon\mathcal{U}\to\mathcal{X}$ is a quasi-compact open immersion. Set $Z:=|\mathcal{X}|\setminus|\mathcal{U}|$. There exists a recollement
$D_{\operatorname{qc}}(\mathcal{U})$$D_{\operatorname{qc}}(\mathcal{X})$$D_{\operatorname{qc},Z}(\mathcal{X})$$\scriptstyle{\mathbf{R}j_{\ast}}$$\scriptstyle{\mathbf{L}j^{\ast}}$$\scriptstyle{j^{\times}}$$\scriptstyle{i^{!}}$$\scriptstyle{i_{\ast}}$$\scriptstyle{i^{\ast}}$
where $i_{\ast}$ is the natural inclusion and $j^{\times}$ the right adjoint of $\mathbf{R}j_{\ast}$.
Proof.
First, we spell out the existence of the needed functors. Recall that the right adjoint $j^{\times}$ of $\mathbf{R}j_{\ast}$ exists because $j$ is concentrated (see [HR17a, Theorem 4.14(1)]). By [HR17b, Example 1.2], there is a Verdier localization
$D_{\operatorname{qc},Z}(\mathcal{X})\xrightarrow{i_{\ast}}D_{\operatorname{qc}}(\mathcal{X})\xrightarrow{\mathbf{L}j^{\ast}}D_{\operatorname{qc}}(\mathcal{U}).$
Note that $\mathbf{R}j_{\ast}$ is right adjoint to $\mathbf{L}j^{\ast}$ on $D_{\operatorname{qc}}$ (see [HR17a, §1.3]). So, by [GP18, Lemma 2.2(ii)], $i_{\ast}$ must admit a right adjoint, which we denote by $i^{!}$ (loc. cit. pulls from [CPS88b, Theorem 1.1] and [CPS88a, Theorem 2.1]). It follows that
$D_{\operatorname{qc},Z}(\mathcal{X})\xleftarrow{i^{!}}D_{\operatorname{qc}}(\mathcal{X})\xleftarrow{\mathbf{R}j_{\ast}}D_{\operatorname{qc}}(\mathcal{U})$
is a Verdier localization sequence. However, being that $j^{\times}$ is the right adjoint of $\mathbf{R}j_{\ast}$ on $D_{\operatorname{qc}}$, we can apply [GP18, Lemma 2.2(ii)] once more to see that $i^{!}$ admits a right adjoint as well, which we denote by $i^{\ast}$. Tying things together, we have the required data for a recollement; i.e. a Verdier localization sequence which is a localization and colocalization sequence in the sense of [KRA10, §4]. ∎
Remark 3.2.
Consider the recollement in Proposition˜3.1. In this case, $i^{!}$ preserves small coproducts, and so, $i_{\ast}$ preserves compact objects (see e.g. the proof of $\implies$ in [NEE96, Theorem 5.1]; which this fact does not require compact generation). Then, from [HR17a, Lemma 4.4(1)], it follows that any compact object of $D_{\operatorname{qc},Z}(\mathcal{X})$ must belong to $\operatorname{Perf}(\mathcal{X})$.
The following is recorded for interest sake.
Corollary 3.3.
Let $\mathcal{X}$ be a quasi-compact quasi-separated algebraic stack. Suppose $j\colon\mathcal{U}\to\mathcal{X}$ is a quasi-compact open immersion. Set $Z:=|\mathcal{X}|\setminus|\mathcal{U}|$. For any closed $W\subseteq|\mathcal{X}|$ such that $D_{\operatorname{qc},|\mathcal{U}|\cap W}(\mathcal{U})$ is compactly generated, there exists a recollement
$D_{\operatorname{qc},|\mathcal{U}|\cap W}(\mathcal{U})$$D_{\operatorname{qc},W}(\mathcal{X})$$D_{\operatorname{qc},W\cap Z}(\mathcal{X})$$\scriptstyle{\mathbf{R}j_{\ast}}$$\scriptstyle{\mathbf{L}j^{\ast}}$$\scriptstyle{j^{!}}$$\scriptstyle{i^{!}}$$\scriptstyle{i_{\ast}}$$\scriptstyle{i^{\ast}}$
where $i_{\ast}$ is the natural inclusion.
Proof.
This mimics the proof of Proposition˜3.1 but a few details need to be spelled out. By Lemma˜2.1, the restriction of $\mathbf{L}j^{\ast}$ to $D_{\operatorname{qc},W}(\mathcal{X})$ factors through $D_{\operatorname{qc},|\mathcal{U}|\cap W}(\mathcal{U})$. Moreover, the restriction of $\mathbf{R}j_{\ast}$ to $D_{\operatorname{qc},|\mathcal{U}|\cap W}(\mathcal{U})$ factors through $D_{\operatorname{qc},W}(\mathcal{X})$, which can be verified using [HR17a, Theorem 2.6(2)] and checking smooth locally. Hence, the adjoint pair $\mathbf{L}j^{\ast}$ and $\mathbf{R}j_{\ast}$ restricts to $D_{\operatorname{qc},W}(\mathcal{X})$ and $D_{\operatorname{qc},|\mathcal{U}|\cap W}(\mathcal{U})$. Furthermore, there is a Verdier localization
$D_{\operatorname{qc},W\cap Z}(\mathcal{X})\xrightarrow{i_{\ast}}D_{\operatorname{qc},W}(\mathcal{X})\xrightarrow{\mathbf{L}j^{\ast}}D_{\operatorname{qc},|\mathcal{U}|\cap W}(\mathcal{U}).$
Indeed, the counit of $\mathbf{L}j^{\ast}$ and $\mathbf{R}j_{\ast}$ is a natural isomorphism on $D_{\operatorname{qc}}$, and so the restriction $\mathbf{L}j^{\ast}\colon D_{\operatorname{qc},W}(\mathcal{X})\to D_{\operatorname{qc},|\mathcal{U}|\cap W}(\mathcal{U})$ is essentially surjective. To see that fully faithfulness holds, use e.g. [GZ67, §I. Proposition 1.3] and [KRA10, Lemma 4.3.1]). Note that restriction of $\mathbf{R}j_{\ast}$ is right adjoint to the restriction of $\mathbf{L}j^{\ast}$. So, by [GP18, Lemma 2.2(ii)], $i_{\ast}$ must admit a right adjoint as well, which we denote by $i^{!}$ (loc. cit. pulls from [CPS88b, Theorem 1.1] and [CPS88a, Theorem 2.1]). It follows that
$D_{\operatorname{qc},Z,\cap W}(\mathcal{X})\xleftarrow{i^{!}}D_{\operatorname{qc},W}(\mathcal{X})\xleftarrow{\mathbf{R}j_{\ast}}D_{\operatorname{qc},|\mathcal{U}|\cap W}(\mathcal{U})$
is a Verdier localization sequence. However, $D_{\operatorname{qc},|\mathcal{U}|\cap W}(\mathcal{U})$ is compactly generated and the restriction of $\mathbf{R}j_{\ast}$ preserves small coproducts. Hence, [BDS16, Corollary 2.3] tells us the restriction of $\mathbf{R}j_{\ast}$ admits a right adjoint. Now, we can apply [GP18, Lemma 2.2(ii)] once more to see that $i^{!}$ admits a right adjoint as well, which we denote by $i^{\ast}$. Tying things together, we have the required data for a recollement; i.e. a Verdier localization sequence which is a localization and colocalization sequence in the sense of [KRA10, §4]. ∎
Remark 3.4.
In Corollary˜3.3, the restriction of $\mathbf{R}j_{\ast}$ to $D_{\operatorname{qc},|\mathcal{U}|\cap W}(\mathcal{U})\to D_{\operatorname{qc},W}(\mathcal{X})$ admits a right adjoint. However, it is not clear to us whether it is the restriction of the right adjoint $j^{\times}$ of $\mathbf{R}j_{\ast}$ on $D_{\operatorname{qc}}$.
3.3. Generating
We record a few useful lemmas allowing one to glue generators along a recollement. Recall that a collection $\mathcal{B}$ in a category $\mathcal{C}$ is said to generate $\mathcal{C}$ if for any nonzero $E\in\mathcal{C}$ there is an $n\in\mathbb{Z}$ and $B\in\mathcal{B}$ such that $\operatorname{Hom}(B[n],E)\not\cong 0$.
Lemma 3.5.
Let $F\colon\mathcal{T}\leftrightarrows\mathcal{S}\colon G$ be an adjoint pair of exact functors between triangulated categories admitting small coproducts. Suppose $\mathcal{T}$ is generated by a collection $\mathcal{B}$. Then $G$ is conservative (i.e. $G(A)\cong 0\implies A\cong 0$) if, and only if, $F(\mathcal{B})$ generates $\mathcal{S}$. In such a case, if $\mathcal{T}$ is generated by a set of cardinality $\leq\beta$ for some cardinal $\beta$, then $\mathcal{B}$ satisfies the same condition.
Proof.
This is known to a few but we spell it out for convienence. Let $E\in\mathcal{B}$ satisfy $\operatorname{Hom}(F(B),E[n])\cong 0$ for all $B\in\mathcal{B}$ and $n\in\mathbb{Z}$. From adjunction, we know that $\operatorname{Hom}(B,G(E)[n])\cong 0$. As $\mathcal{B}$ generates $\mathcal{T}$, it follows that $G(E)\cong 0$. However, $G$ being conservative implies $E\cong 0$. So, $F(\mathcal{B})$ generates $\mathcal{S}$.
Conversely, assume that $F(\mathcal{B})$ generates $\mathcal{S}$. Let $E\in\mathcal{S}$ satisfy $G(E)\cong 0$. By adjunction, it follows that $0\cong\operatorname{Hom}(F(B),E[n])$ for all $B\in\mathcal{B}$ and $n\in\mathbb{Z}$. Yet, the assumption implies $E\cong 0$. Hence, $G$ must be conservative.
That the last claim holds follows from the proof above. ∎
Corollary 3.6.
Let $F\colon\mathcal{T}\leftrightarrows\mathcal{S}\colon G$ be an adjoint pair of exact functors between triangulated categories admitting small coproducts. Suppose $\mathcal{T}$ is compactly generated by a collection $\mathcal{B}$ and $G$ commutes with small coproducts. Then $G$ is conservative if, and only if, $F(\mathcal{B})$ compactly generates $\mathcal{S}$. In such a case, if $\mathcal{T}$ is compactly generated by a set of cardinality $\leq\beta$ for some cardinal $\beta$, then $\mathcal{B}$ satisfies the same condition.
Proof.
That $G$ commutes with small coproducts ensures that $F(\mathcal{T}^{c})\subseteq\mathcal{S}^{c}$ (see e.g. the proof of $\implies$ in [NEE96, Theorem 5.1]; which this fact does not require $\mathcal{T}$ to be compactly generated). Consequently, the desired claim follows from Lemma˜3.5. ∎
Proposition 3.7.
Consider a recollement as in Diagram˜3.1. Let $\mathcal{G}$ be a subcategory of $\mathcal{K}$ such that $I_{\lambda}(\mathcal{G})$ generates $\mathcal{T}$. If $\mathcal{G}^{\prime}$ is a subcategory which generates $\mathcal{D}$, then the collection of $G\oplus Q_{\lambda}(G^{\prime})$ where $G\in\mathcal{G}$ and $G^{\prime}\in\mathcal{G}^{\prime}$ generates $\mathcal{K}$.
Proof.
Let $E\in\mathcal{K}$ satisfy $\operatorname{Hom}((G\oplus Q_{\lambda}(G^{\prime}))[n],E)=0$ for all $n\in\mathbb{Z}$, $G\in\mathcal{G}$, and $G^{\prime}\in\mathcal{G}^{\prime}$. From the recollement above, there is a distinguished triangle
$(Q_{\lambda}\circ Q)(E)\to E\to(I\circ I_{\lambda})(E)\to(Q_{\lambda}\circ Q)(E)[1].$
Here, in such a case, we have that $\operatorname{Hom}(Q_{\lambda}(G^{\prime})[n],E)\cong 0$ for all $n\in\mathbb{Z}$. So, adjunction tells us $\operatorname{Hom}(G^{\prime}[n],Q(E))\cong 0$ for all $n\in\mathbb{Z}$. However, $\mathcal{G}^{\prime}$ generates $\mathcal{D}$, so $Q(E)\cong 0$. This implies the morphism $E\to(I\circ I_{\lambda})(E)$ from the distinguished triangle above is an isomorphism. Once more, from adjunction, we then have that $\operatorname{Hom}(G[n],(I\circ I_{\lambda})(E))=0$ for all $n\in\mathbb{Z}$. So, $\operatorname{Hom}(I_{\lambda}(G)[n],I_{\lambda}(E))=0$ via adjunction for any such $n$. However, $I_{\lambda}(\mathcal{G})$ generates $\mathcal{T}$, so $I_{\lambda}(E)$. Consequently, $E\cong 0$, which completes the proof. ∎
4. Proofs
This section containes our main results. To start, we prove the following. It is useful in later proofs where we leverage a duality trick.
Lemma 4.1.
Let $f\colon\mathcal{Y}\to\mathcal{X}$ be a finite morphism of Noetherian algebraic stacks. Then the adjoint pair $\mathbf{R}f_{\ast}$ and $f^{\times}$ restricts to $D_{\operatorname{qc}}(\mathcal{Y})$ and $D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X})$. Additionally, if $D_{\operatorname{qc}}(\mathcal{Y})$ is generated by some $\mathcal{B}\subseteq D^{b}_{\operatorname{coh}}(\mathcal{X})$ and $\mathbf{R}f_{\ast}\mathcal{O}_{\mathcal{Y}}\in\operatorname{Perf}(\mathcal{X})$, then $\mathbf{R}f_{\ast}\mathcal{B}$ generates $D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X})$.
Proof.
To start, we show the claim regarding restricting the adjunction. Note that a finite morphism is a closed mapping on the underlying topological spaces. Recall, by [HR17a, Theorem 4.14(1)], the right adjoint of $\mathbf{R}f_{\ast}$ on $D_{\operatorname{qc}}$ exists because $f$ is concentrated. It is easy to see that $f^{\times}(D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X}))\subseteq D_{\operatorname{qc}}(\mathcal{Y})$. Also, by Lemma˜2.2, we know that $\mathbf{R}f_{\ast}(D_{\operatorname{qc}}(\mathcal{Y}))\subseteq D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X})$ because $f$ is finite.
Lastly, we show that $\mathbf{R}f_{\ast}\mathcal{B}$ generates $D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X})$. By Lemma˜3.5, it suffices to show that $f^{\times}\colon D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X})\to D_{\operatorname{qc}}(\mathcal{Y})$ is conservative. So, let $E\in D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X})$ satisfy $f^{\times}E\cong 0$. By [HR17a, Theorem 4.14(2)], it follows that
$0\cong\mathbf{R}f_{\ast}f^{\times}E\cong\operatorname{\mathbf{R}\mathcal{H}\!\mathit{om}}(\mathbf{R}f_{\ast}\mathcal{O}_{\mathcal{Y}},E).$
However, $\mathbf{R}f_{\ast}\mathcal{O}_{\mathcal{Y}}\in\operatorname{Perf}(\mathcal{X})\cap D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X})$, and so, [HR17a, Lemma 4.9] ensures that $E\cong 0$. Hence, we see that the restriction of $f^{\times}$ on $D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X})$ is conservative. ∎
Remark 4.2.
In Lemma˜4.1, [HR17a, Theorem 4.14(4)] tells us $f^{\times}$ preserves small coproducts. So, Corollary˜3.6 tells us $\mathbf{R}f_{\ast}\mathcal{B}$ compactly generates $D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X})$ if $\mathcal{B}$ does such for $D_{\operatorname{qc}}(\mathcal{Y})$.
Next, we prove a variation of [HR17a, Theorem 6.6] for the $\beta$-Thomason condition. Although it is used in the proof of Theorem˜1.1, it is possible to replace references to it with [HR17a, Theorem 6.6]. So, we record it for interest sake.
Proposition 4.3.
Let $f\colon\mathcal{Y}\to\mathcal{X}$ be a finite, flat, surjective, morphism between Noetherian algebraic stacks. If $\mathcal{Y}$ satisfies the $\beta$-Thomason condition, then so does $\mathcal{X}$.
Proof.
Since $f\colon\mathcal{Y}\to\mathcal{X}$ is a finite, flat, surjective morphism of Noetherian algebraic stacks, it follows that $\mathbf{R}f_{\ast}\mathcal{O}_{\mathcal{Y}}\in\operatorname{Perf}(\mathcal{X})$. Denote by $f^{\times}$ the right adjoint of $\mathbf{R}f_{\ast}$ on $D_{\operatorname{qc}}$. In our case, [HR17a, Theorem 4.14] tells us that $f^{\times}$ preserves small coproducts and is conservative. By the hypothesis, we know that there exists a collection $\mathcal{B}$ (of some cardinality $\leq\beta$) which compactly generates $D_{\operatorname{qc}}(\mathcal{Y})$.
Choose a closed $Z\subseteq|\mathcal{X}|$. The hypothesis on $\mathcal{Y}$ tells us there is a $P\in\operatorname{Perf}(\mathcal{Y})$ such that $\operatorname{supp}(P)=f^{-1}(Z)$. So, appealing to [HR17a, Lemma 4.10(2)], we know that $D_{\operatorname{qc},f^{-1}(Z)}(\mathcal{Y})$ is compactly generated by objects of the form $B\otimes^{\mathbf{L}}P$ with $B\in\mathcal{B}$.
We claim that the collection of $\mathbf{R}f_{\ast}(B\otimes^{\mathbf{L}}P)$ with $B\in\mathcal{B}$ compactly generates $D_{\operatorname{qc},Z}(\mathcal{X})$. Towards that end, we show that the adjunction $\mathbf{R}f_{\ast}\dashv f^{\times}$ restricts to an adjoint pair between $D_{\operatorname{qc},f^{-1}(Z)}(\mathcal{Y})$ and $D_{\operatorname{qc},Z}(\mathcal{X})$. By [HR17a, Theorem 4.14(4)], we know that $f^{\times}\mathcal{O}_{\mathcal{X}}\otimes^{\mathbf{L}}\mathbf{L}f^{\ast}E\cong f^{\times}E$ for all $E\in D_{\operatorname{qc},Z}(\mathcal{X})$. This ensures that $\operatorname{supp}(f^{\times}(E))\subseteq\operatorname{supp}(\mathbf{L}f^{\ast}E)$. Moreover, Lemma˜2.1 promises that $\operatorname{supp}(\mathbf{L}f^{\ast}E)\subseteq f^{-1}(Z)$, and so $f^{\times}D_{\operatorname{qc},Z}(\mathcal{X})\subseteq D_{\operatorname{qc},f^{-1}(Z)}(\mathcal{Y})$.
Next, we check that $\mathbf{R}f_{\ast}D_{\operatorname{qc},f^{-1}(Z)}(\mathcal{Y})\subseteq D_{\operatorname{qc},Z}(\mathcal{X})$. Recall that $D_{\operatorname{qc},f^{-1}(Z)}(\mathcal{Y})$ is compactly generated by the collection $B\otimes^{\mathbf{L}}P$ with $B\in\mathcal{B}$. Hence, every object of $D_{\operatorname{qc},f^{-1}(Z)}(\mathcal{Y})$ is a homotopy colimit of iterated extensions of small coproducts of shifts of the collection $B\otimes^{\mathbf{L}}P$ (see [STA26, Tag 09SN]). So, this can be checked by showing $\mathbf{R}f_{\ast}(B\otimes^{\mathbf{L}}P)\in D_{\operatorname{qc},Z}(\mathcal{X})$. However, this can be seen by reducing to schemes (see e.g. [GW23, Remark 23.46(2)]).
Now, we finish the proof. As we have shown the adjunction restricts, we can make use of the fact that $f^{\times}$ is conservative by [HR17a, Theorem 4.14(4)]. Consequently, Corollary˜3.6 implies the collection of $\mathbf{R}f_{\ast}(B\otimes^{\mathbf{L}}P)$ with $B\in\mathcal{B}$ compactly generates $D_{\operatorname{qc},Z}(\mathcal{X})$. This tells us every object of $D_{\operatorname{qc},Z}(\mathcal{X})$ is a homotopy colimit of iterated extensions of small coproducts of shifts of the collection $\mathbf{R}f_{\ast}(B\otimes^{\mathbf{L}}P)$ with $B\in\mathcal{B}$ (see [STA26, Tag 09SN]). Hence, if the support of every $\mathbf{R}f_{\ast}(B\otimes^{\mathbf{L}}P)$ is properly contained in $Z$, then so must the support of every object in $D_{\operatorname{qc},Z}(\mathcal{X})$, which is absurd. Therefore, there must exist some $B\in\mathcal{B}$ such that $\operatorname{supp}(\mathbf{R}f_{\ast}(B\otimes^{\mathbf{L}}P))=Z$. ∎
Secondly, we need a version of [NEE23, Observation 5.6] which is applicable to quasi-affine monomorphisms, which appears in the proof of Theorem˜1.1.
Lemma 4.4.
Let $t\colon\mathcal{Y}\to\mathcal{X}$ be a quasi-affine morphism of quasi-compact quasi-separated algebraic stacks. Suppose $t$ is injective on the underlying topological spaces. If $\mathcal{X}$ satisfies the $\beta$-Thomason condition, then so does $\mathcal{Y}$.
Proof.
As $t$ is quasi-affine, it must be concentrated. So, [HR17a, Lemma 8.2] guarantees that $D_{\operatorname{qc}}(\mathcal{Y})$ is compactly generated by a collection of at most $\leq\beta$ objects. Next, let $Z$ be a closed subset of $|\mathcal{Y}|$. Set $Z^{\prime}$ to be the closure of $t(Z)$ in $|\mathcal{X}|$. Then $t^{-1}(Z^{\prime})=Z$ because $t$ is injective on the underlying topological spaces. Moreover, the hypothesis on $\mathcal{X}$ allows us to find a $P\in\operatorname{Perf}(\mathcal{X})$ such that $\operatorname{supp}(P)=Z^{\prime}$. However, [HR17a, Lemma 4.8(2)] tells us $t^{-1}(\operatorname{supp}(P))=\operatorname{supp}(\mathbf{L}t^{\ast}P)$, which completes the proof because $\mathbf{L}t^{\ast}P\in\operatorname{Perf}(\mathcal{Y})$. ∎
Now, we prove the main result.
Proof of Theorem˜1.1.
The proof proceeds by an inductive argument. However, to set the stage, we start with a construction needed for such an argument. By [HR18, Theorem 4.1], there exist morphisms of algebraic stacks $V\xrightarrow{p}\mathcal{Y}\xrightarrow{f}\mathcal{X}$ satisfying
  • $V$ is an affine scheme
  • $p$ is finite, flat, surjective, and of finite presentation
  • $f$ is a Nisnevich covering (see [HR18, Definition 3.1]) of finite presentation with separated diagonal (in fact, $f$ is étale and surjective).
Moreover, from [HR18, Proposition 3.1], there exists a monomorphic splitting sequence for $f$ in the sense of loc. cit. Specfically, we have a sequence of open immersions
$\emptyset=:\mathcal{X}_{0}\xrightarrow{j_{0}}\mathcal{X}_{1}\xrightarrow{j_{1}}\cdots\xrightarrow{j_{n-2}}\mathcal{X}_{n-1}\xrightarrow{j_{n-1}}\mathcal{X}_{n}=:\mathcal{X}$
such that the base change of $f$ to $|\mathcal{X}_{c}|\setminus|\mathcal{X}_{c-1}|$ (when given the reduced closed substack structure) admits a monomorphic section for each $c=1,\ldots,n$.
Next, before going to the induction argument, we make an observation for components of the construction above. Fix $c=1,\ldots,n$. Consider the fibered square
$\mathcal{Y}_{c}$$(|\mathcal{X}_{c}|\setminus|\mathcal{X}_{c-1}|)_{red}$$\mathcal{Y}$$\mathcal{X}$$\scriptstyle{f_{c}}$$\scriptstyle{h_{c}}$$\scriptstyle{j^{\prime}_{c}\circ t_{c}}$$\scriptstyle{f}$
obtained by base change of $f$ along $j^{\prime}_{c}\circ t_{c}$ where $j^{\prime}_{c}$ is the associated open immersion and $t_{c}$ the associated closed immersion. Choose a monomorphic section $g_{c}\colon(|\mathcal{X}_{c}|\setminus|\mathcal{X}_{c-1}|)_{red}\to\mathcal{Y}_{c}$. This gives us a commutative diagram
[…]
Base change tells us that $f_{c}$ is étale and of finite presentation. Hence, $g_{c}$ must be étale (see e.g. [STA26, Tag 0CIR]). Furthermore, since $g_{c}$ is monomorphic, it is separated, of finite type, and representable by algebraic spaces (see e.g. [STA26, Tag 06MY & Tag 050J]). Yet, each algebraic stack in our proof is Noetherian, so $g_{c}$ must be of finite presentation (see e.g. [STA26, Tag 0DQJ]). It follows that $g_{c}$ must be quasi-affine (use e.g. [OS03, Proposition 3.1]). There is a fibered square,
[…]
Again, from base change, we know that $h_{c}$, and hence $h^{\prime}_{c}$, is an immersion. Since any such morphism is quasi-affine, it follows that $V_{c}$ is a quasi-affine scheme. Hence, by Proposition˜4.3, we know that $\mathcal{Y}_{c}$ must be $1$-Thomason because $p_{c}$ is a finite, flat, and surjective morphism of finite presentation from a quasi-affine scheme. Thus, from Lemma˜4.4, we know that each $(|\mathcal{X}_{c}|\setminus|\mathcal{X}_{c-1}|)_{red}$ must be $1$-Thomason because $g_{c}$ is a quasi-affine monomorphism morphism (see e.g. [STA26, Tag 0500]). Consequently, as each $\mathcal{X}_{c}$ is regular, Lemma˜4.1 tells us that $D_{\operatorname{qc},Z_{c}}(\mathcal{X_{c}})$ is singly compactly generated where $Z_{c}:=t_{c}(|(|\mathcal{X}_{c}|\setminus|\mathcal{X}_{c-1}|)_{red}|)$. So, we may find a $P_{c}\in\operatorname{Perf}(\mathcal{X}_{c})$ which compactly generates $D_{\operatorname{qc},Z_{c}}(\mathcal{X_{c}})$.
Now, we prove the desired claim by induction on $n$. Specifically, we are inducting the length of the monomorphic splitting sequence of $f$. There is nothing to check if $n=0$. Moreover, the observation above addresses the base case $n=1$. So, we may assume that $n\geq 2$. By [HLL+25, Proposition B.1], there is a Verdier localization sequence
$D^{b}_{\operatorname{coh},Z_{c}}(\mathcal{X}_{c})\xrightarrow{(i_{c})_{\ast}}D^{b}_{\operatorname{coh}}(\mathcal{X}_{c})\xrightarrow{\mathbf{L}j_{c-1}^{\ast}}D^{b}_{\operatorname{coh}}(\mathcal{X}_{c-1})$
where $(i_{c})_{\ast}$ is the natural inclusion. However, [DLM+25a, Theorem 3.7] tells us that $\operatorname{Perf}=D^{b}_{\operatorname{coh}}$ in each case. This allows us to find a $G_{c}\in\operatorname{Perf}(\mathcal{X}_{c})$ such that $\mathbf{L}j_{c-1}^{\ast}G_{c}$ generates $D_{\operatorname{qc}}(\mathcal{X}_{c-1})$ and belongs to $\operatorname{Perf}(\mathcal{X}_{c-1})$. Consider the recollement obtained in Proposition˜3.1,
$D_{\operatorname{qc}}(\mathcal{X}_{c-1})$$D_{\operatorname{qc}}(\mathcal{X}_{c})$$D_{\operatorname{qc},Z_{c}}(\mathcal{X}_{c})$$\scriptstyle{\mathbf{R}(j_{c-1})_{\ast}}$$\scriptstyle{\mathbf{L}j^{\ast}_{c-1}}$$\scriptstyle{j^{\times}_{c}}$$\scriptstyle{i^{!}_{c}}$$\scriptstyle{(i_{c})_{\ast}}$$\scriptstyle{i_{c}^{\ast}}$
where $(i_{c})_{\ast}$ is the natural inclusion and $j^{\times}_{c}$ the right adjoint of $\mathbf{R}j_{\ast}$. So, in particular, $(i_{c})_{\ast}P_{c}=P_{c}$. Then, from Proposition˜3.7, we know that $G_{c}\oplus P_{c}$ generates $D_{\operatorname{qc}}(\mathcal{X}_{c})$. Consequently, this establishes the induction step. Since $f$ always has monomorphic splitting sequence of finite length, we complete the proof. ∎
Remark 4.5.
In the proof of Theorem˜1.1, if $G_{c}$ can be chosen to be compact, then it follows that $D_{\operatorname{qc}}(\mathcal{X}_{c})$ is singly compactly generated. However, it is not clear to us whether this is always possible.
Lemma 4.6.
Let $\mathcal{X}$ be a quasi-compact quasi-separated algebraic stack. Suppose $\mathcal{S}\subseteq\operatorname{Perf}(\mathcal{X})$ generates $D_{\operatorname{qc}}(\mathcal{X})$. Consider a quasi-compact open immersion $j\colon\mathcal{U}\to\mathcal{X}$ with complement $Z:=|\mathcal{X}|\setminus|\mathcal{U}|$. If there is a $Q\in\operatorname{Perf}(\mathcal{X})$ such that $\operatorname{supp}(Q)=Z$, then $D_{\operatorname{qc},Z}(\mathcal{X})$ is generated by the collection $S\otimes^{\mathbf{L}}Q$ where $S\in\mathcal{S}$.
Proof.
This mimics the proof of [HR17a, Lemma 4.10(2)], but without compactness. Suppose $E\in D_{\operatorname{qc},Z}(\mathcal{X})$ satisfies $\operatorname{Hom}(S\otimes^{\mathbf{L}}Q[n],E)\cong 0$ for all $S\in\mathcal{S}$ and $n\in\mathbb{Z}$. By adjunction, it follows that $\operatorname{Hom}(S[n],\operatorname{\mathbf{R}\mathcal{H}\!\mathit{om}}(Q,E))\cong 0$. However, $\mathcal{S}$ generates $D_{\operatorname{qc}}(\mathcal{X})$, and so $\operatorname{\mathbf{R}\mathcal{H}\!\mathit{om}}(Q,E)\cong 0$. Since $\operatorname{supp}(Q)=Z$, it follows from [HR17a, Lemma 4.9] that $E\cong 0$, which completes the proof. ∎
Corollary 4.7.
Let $\mathcal{X}$ be a quasi-compact quasi-separated regular algebraic stack with quasi-finite diagonal. Then $D_{\operatorname{qc},Z}(\mathcal{X})$ is generated by a single perfect complex for all $Z=|\mathcal{X}|\setminus|\mathcal{U}|$ with $\mathcal{U}\to\mathcal{X}$ a quasi-compact open immersion.
Proof.
Since $\mathcal{X}$ is regular and Noetherian, we know that $\operatorname{Perf}(\mathcal{X})=D^{b}_{\operatorname{coh}}(\mathcal{X})$ (see e.g. [DLM+25a, Theorem 3.7]). Denote by $i\colon\mathcal{Z}\to\mathcal{X}$ for the closed immersion from the reduced induced closed substack structure on $Z$. Hence, $\mathbf{R}i_{\ast}\mathcal{O}_{\mathcal{Z}}\in\operatorname{Perf}(\mathcal{X})$ with $\operatorname{supp}(\mathbf{R}i_{\ast}\mathcal{O}_{\mathcal{Z}})=Z$. Consequently, the desired claim follows from Theorem˜1.1 and Lemma˜4.6. ∎
Proof of Corollary˜1.2.
If $\mathcal{X}$ is concentrated, then $\operatorname{Perf}(\mathcal{X})=D_{\operatorname{qc}}(\mathcal{X})^{c}$. So, the desired claim follows from Corollary˜4.7 with $Z=|\mathcal{X}|$. In fact, this shows that $\mathcal{X}$ satisfies the $1$-Thomason condition, see Remark˜3.2. ∎